OpenStudy (anonymous):

How do I solve this system by substituting? t = 2r + 3 5r - 4t = 6

hero (hero):

\[t = 2r + 3\] \[t = \frac{6-5r}{-4} \] Set t = t: \[2r + 3 = \frac{6-5r}{-4} \] -4(2r+3) = 6 - 5r -8r-12 = 6 - 5r 3r = -18 r = -6 t = -9